Find the distance between the two points
(3,7) and (-12,7)

Answer by venugopalramana(3286)

Find the distance between the two points
(x1,y1)=(3,7) and (x2,y2)=(-12,7)
distance formula is
d=sqrt.of[(x2-x1)^2+(y2-y1)^2]=sqrt.of[(-12-3)^2+(7-7)^2]=sqrt.of[15^2]=15